OpenStudy (anonymous):

An equation for line k is y=-3/8x+5. An equation for line p is 8x + 3y = 10. Which best describes the relationship between line k and line p? A. k || p B. k ⊥ p C. They are neither perpendicular nor parallel. D. They are the same line.

hartnn (hartnn):

isolate 'y' from her 8x + 3y = 10 in other words, express it in y=mx+c form

OpenStudy (anonymous):

3y = -8x +10

OpenStudy (anonymous):

is this correct?

hartnn (hartnn):

correct so far, then ?

OpenStudy (anonymous):

find the slopes...

hartnn (hartnn):

if you can find the slopes, and if m1=m2, lines are parallel, m1*m2 = -1, lines are perpendicular, else neither

hartnn (hartnn):

if the equation comes out to be the same entirely, then they are same line.

OpenStudy (anonymous):

So the answer is D ?

hartnn (hartnn):

are the equations exactly same ?

OpenStudy (anonymous):

not exactly

hartnn (hartnn):

then its not D

OpenStudy (anonymous):

They have the same slopes, so are they parallel then?

hartnn (hartnn):

are you sure they have same slopes ? what are the slopes according to you ?

OpenStudy (anonymous):

No I am sorry my mistake, they dont

hartnn (hartnn):

what about m1*m2 =-1 is this true ?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

so is the correct answer B?

hartnn (hartnn):

how ? can you tell me what m1=...? m2=.... ?

OpenStudy (anonymous):

I am sorry. I am confused

hartnn (hartnn):

you got -3/8 and -8/3 , right ? now, -3/8 * -8/3 = 1 and not -1 so they are not perpendicular also. hence, its 'They are neither perpendicular nor parallel.'

OpenStudy (anonymous):

Ooh so the right answer is C?

hartnn (hartnn):

yes.

OpenStudy (anonymous):

Thank you very much :(

hartnn (hartnn):

why thanks with a sad face ?

OpenStudy (anonymous):

Thanks for clearing it up I kinda have a really hard time solving these problems:(

hartnn (hartnn):

practice makes a man perfect ! (and women also :P) so, keep practicing and you'll be good at it :)

OpenStudy (anonymous):

Thanks for the support!!!! I REALLY appreciate it!!! :)